OpenStudy (anonymous):

Given the circle with radius 12, what is the probability of choosing a point inside the triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

1) Fist of all find area of circle, it will be 3.14 * 12 * 12 = 450 2) So the area of triangle will be 1/2 * 12 * 12 * sin 120 (0.86) = 61 3) Probability = 61 / 450 = 0.13 (Approximate)